# Forced Symmetric Formation Control

### v1

@@ IV-C Symmetry forced formations with centroid consensus
The $\tau(\Gamma)$-symmetric frameworks by definition have the point-group symmetries defined with respect to some fixed inertial point (the origin). As seen in Figure 6(d), for agent initial conditions that are far from the origin, the control strategy in (25) will drive the agent to the correct formation, but with respect to the origin. In fact, it would be more desirable for the agents to be able to arrange themselves to the correct symmetric formation with respect to any point. This is further a necessary requirement if we want to include formation maneuvering as well.
@@
We now are able to use the virtual state $r_{i}(t)$ for each agent to effectively shift the origin in the control (25).
@@
$\begin{bmatrix}\dot{p}_{0}(t)\\ \dot{p}_{f}(t)\end{bmatrix}$ $=\begin{bmatrix}-\mathcal{O}^{T}(\mathcal{G}_{0},c_{0}(t))\bigg{(}\mathcal{O}(\mathcal{G}_{0},c_{0}(t))c_{0}(t)-\mathrm{\bf d}_{0}^{2}\bigg{)}\\ 0\end{bmatrix}$ (52)
@@
The combined dynamics (46) and (52) are in a cascade form. Following the same approach as before, we define the error signals
@@
$\begin{bmatrix}\dot{\bar{\sigma}}(t)\\ \dot{\bar{q}}(t)\end{bmatrix}$ $=-\begin{bmatrix}\mathcal{O}(\mathcal{G}_{0},c_{0}(t))\mathcal{O}^{T}(\mathcal{G}_{0},c_{0}(t))&\mathcal{O}(\mathcal{G}_{0},c_{0}(t))\bar{E}_{0}(\Gamma)\\ \bar{E}_{0}^{T}(\Gamma)\mathcal{O}^{T}(\mathcal{G}_{0},c_{0}(t))&\bar{E}^{T}(\Gamma)\bar{E}(\Gamma)\end{bmatrix}\begin{bmatrix}{\bar{\sigma}}(t)\\ {\bar{q}}(t)\end{bmatrix}$ (62)
@@
To assist in our analysis of the cascade system (62) and (46), we perform a simple change of coordinates. Let $\hat{p}(t)=p(t)-{\mathbf{r}}^{\star}$, $\hat{r}(t)=r(t)-{\mathbf{r}}^{\star}$. With this definition, note that $\hat{c}(t)=\hat{p}(t)-\hat{r}(t)=c(t)$, similarly for the error signals $\bar{\sigma}(t)$ and $\bar{q}(t)$ (i.e., $\hat{\sigma}(t)=\bar{\sigma}(t)$ and $\hat{q}(t)=\bar{q}(t)$). With this definition, the shifted cascade system can be represented as
$\begin{bmatrix}\dot{\hat{\sigma}}(t)\\ \dot{\hat{q}}(t)\end{bmatrix}$ $=-\begin{bmatrix}\mathcal{O}(\mathcal{G}_{0},\hat{c}_{0}(t))\mathcal{O}^{T}(\mathcal{G}_{0},\hat{c}_{0}(t))&\mathcal{O}(\mathcal{G}_{0},\hat{c}_{0}(t))\bar{E}_{0}(\Gamma)\\ \bar{E}_{0}^{T}(\Gamma)\mathcal{O}^{T}(\mathcal{G}_{0},\hat{c}_{0}(t))&\bar{E}^{T}(\Gamma)\bar{E}(\Gamma)\end{bmatrix}\begin{bmatrix}{\hat{\sigma}}(t)\\ {\hat{q}}(t)\end{bmatrix}$ (73)
@@
First we establish that (77) is locally input-to-state stable while interpreting the signal $\hat{r}(t)$ as the input. When $\hat{r}(t)=0$ we have that $\hat{c}_{0}(t)=\hat{p}_{0}(t)$ and the dynamics (77) reduce to (35). Using the same arguments as in Theorem 4, this system is locally exponentially stable, which shows the system must be locally input-to-state stable [34]. To establish that the cascade system is exponentially stable, all we need is to recall that (78) is exponentially stable and converges from any initial condition to the origin [33]. It can then be concluded that the cascade system is also locally exponentially stable [35]. ∎
@@
Consider the cascade dynamics (52) and (46) and assume that $(\mathcal{G},p^{\star}-{\mathbf{r}}^{\star})$ is a $\tau(\Gamma)$-symmetric isostatic framework. Then for all initial conditions sufficiently close to $p^{\star}$, and ${\mathbf{r}}^{\star}$ given in (47), $p(t)$ exponentially converges to $p^{\star}$.
@@
We return to the same setup as Example 6 and now implement the centroid consensus version of the dynamics (52). The resulting trajectories can be seen in Figure 6(e). Here we initialize the virtual state $r(t)$ to be the initial condition of the agents, $r(0)=p(0)$. The formation converges to the correct symmetry-forced configuration, but now with respect to the centroid of the initial conditions.
@@ V Concluding Remarks
This work presented a formation control strategy for controlling a network of agents to a formation characterized by its symmetry properties. Leveraging recent results from symmetry-forced rigidity theory, a gradient control strategy was derived with two main components based on a chosen symmetry group of the formation: one maintains distances between edges in the edge orbits, while the other forces agents within the same vertex orbit to a symmetric position. The stability results turn out to be related to properties of the so-called orbit rigidity matrix which is a central object in symmetry-force rigidity theory. While the results in this work focused only on symmetries that are free, extending these ideas to the non-free case should be straightforward and is the topic of future research. Additional research directions include exploring symmetry-preserving motions for formation maneuvering.

### v2

@@ IV-C Symmetry forced formations with centroid consensus
The $\tau(\Gamma)$-symmetric frameworks by definition have the point-group symmetries defined with respect to some fixed inertial point (the origin). As seen in Figure 6(d), for agent initial conditions that are far from the origin, the control strategy in (IV-B) will drive the agent to the correct formation, but with respect to the origin. In fact, it would be more desirable for the agents to be able to arrange themselves to the correct symmetric formation with respect to any point. This is further a necessary requirement if we want to include formation maneuvering as well.
@@
We now are able to use the virtual state $r_{i}(t)$ for each agent to effectively shift the origin in the control (IV-B).
@@
$\begin{bmatrix}\dot{p}_{0}(t)\\ \dot{p}_{f}(t)\end{bmatrix}$ $=\begin{bmatrix}-\mathcal{O}^{T}(\mathcal{G}_{0},c_{0}(t))\bigg{(}\mathcal{O}(\mathcal{G}_{0},c_{0}(t))c_{0}(t)-\mathrm{\bf d}_{0}^{2}\bigg{)}\\ 0\end{bmatrix}$
@@
The combined dynamics (21) and (IV-C) are in a cascade form. Following the same approach as before, we define the error signals
@@
$\begin{bmatrix}\dot{\bar{\sigma}}(t)\\ \dot{\bar{q}}(t)\end{bmatrix}$ $=-\begin{bmatrix}\mathcal{O}(\mathcal{G}_{0},c_{0}(t))\mathcal{O}^{T}(\mathcal{G}_{0},c_{0}(t))&\mathcal{O}(\mathcal{G}_{0},c_{0}(t))\bar{E}_{0}(\Gamma)\\ \bar{E}_{0}^{T}(\Gamma)\mathcal{O}^{T}(\mathcal{G}_{0},c_{0}(t))&\bar{E}^{T}(\Gamma)\bar{E}(\Gamma)\end{bmatrix}\begin{bmatrix}{\bar{\sigma}}(t)\\ {\bar{q}}(t)\end{bmatrix}$
@@
To assist in our analysis of the cascade system (IV-C) and (21), we perform a simple change of coordinates. Let $\hat{p}(t)=p(t)-{\mathbf{r}}^{\star}$, $\hat{r}(t)=r(t)-{\mathbf{r}}^{\star}$. With this definition, note that $\hat{c}(t)=\hat{p}(t)-\hat{r}(t)=c(t)$, similarly for the error signals $\bar{\sigma}(t)$ and $\bar{q}(t)$ (i.e., $\hat{\sigma}(t)=\bar{\sigma}(t)$ and $\hat{q}(t)=\bar{q}(t)$). With this definition, the shifted cascade system can be represented as
$\begin{bmatrix}\dot{\hat{\sigma}}(t)\\ \dot{\hat{q}}(t)\end{bmatrix}$ $=-\begin{bmatrix}\mathcal{O}(\mathcal{G}_{0},\hat{c}_{0}(t))\mathcal{O}^{T}(\mathcal{G}_{0},\hat{c}_{0}(t))&\mathcal{O}(\mathcal{G}_{0},\hat{c}_{0}(t))\bar{E}_{0}(\Gamma)\\ \bar{E}_{0}^{T}(\Gamma)\mathcal{O}^{T}(\mathcal{G}_{0},\hat{c}_{0}(t))&\bar{E}^{T}(\Gamma)\bar{E}(\Gamma)\end{bmatrix}\begin{bmatrix}{\hat{\sigma}}(t)\\ {\hat{q}}(t)\end{bmatrix}$
@@
First we establish that (25) is locally input-to-state stable while interpreting the signal $\hat{r}(t)$ as the input. When $\hat{r}(t)=0$ we have that $\hat{c}_{0}(t)=\hat{p}_{0}(t)$ and the dynamics (25) reduce to (IV-B). Using the same arguments as in Theorem 4, this system is locally exponentially stable, which shows the system must be locally input-to-state stable [32]. To establish that the cascade system is exponentially stable, all we need is to recall that (26) is exponentially stable and converges from any initial condition to the origin [31]. It can then be concluded that the cascade system is also locally exponentially stable [33]. ∎
@@
Consider the cascade dynamics (IV-C) and (21) and assume that $(\mathcal{G},p^{\star}-{\mathbf{r}}^{\star})$ is a $\tau(\Gamma)$-symmetric isostatic framework. Then for all initial conditions sufficiently close to $p^{\star}$, and ${\mathbf{r}}^{\star}$ given in (22), $p(t)$ exponentially converges to $p^{\star}$.
@@
We return to the same setup as Example 6 and now implement the centroid consensus version of the dynamics (IV-C). The resulting trajectories can be seen in Figure 6(e). Here we initialize the virtual state $r(t)$ to be the initial condition of the agents, $r(0)=p(0)$. The formation converges to the correct symmetry-forced configuration, but now with respect to the centroid of the initial conditions.
@@ V Concluding Remarks
This work presented a formation control strategy for controlling a network of agents to a formation characterized by its symmetry properties. Leveraging recent results from symmetry-forced rigidity theory, a gradient control strategy was derived with two main components based on a chosen symmetry group of the formation: one maintains distances between edges in the edge orbits, while the other forces agents within the same vertex orbit to a symmetric position. The stability results turn out to be related to properties of the so-called orbit rigidity matrix which is a central object in symmetry-force rigidity theory. While the results in this work focused only on symmetries that are free, extending these ideas to the non-free case should be straightforward and is the topic of future research.